D2n Run Plan 2020

From HallCWiki
Jump to: navigation, search

Contents

Rough / High-level sketch of d2n kinematic settings

Reduced kinematics list to accommodate limited beam time (dropped 1/4 of kinematics)

      ---------  Running Conditions ----------
E_beam = 10.4 GeV/c     I_beam = 30 uA      P_beam = 0.80

          ---------  HMS ----------
       E'    theta    x        Q2       W
 A    -4.2   13.5     0.207    2.414   3.178 
 B    -4.2   16.4     0.305    3.554   2.993
 C    -4.0   20.0     0.418    5.018   2.806 

          --------- SHMS ----------
       E'    theta    x       Q2      W
 X    -7.5   11.0     0.527   2.866   1.859
 Y    -6.4   14.5     0.565   4.240   2.036
 Z    -5.6   18.0     0.633   5.701   2.046

Run / Shift Schedule

  • We started taking on Swing on Aug 7 (took a few days for beam to stabilize)
  • Beam is scheduled to turn off 8am, Sept 21 (roughly 6 weeks) [1]
  • Hall C Shift Schedule

Data Collection Status

  • Data collection status is best monitored using d2n_status_table on cdaq@cdaql1
[cdaq@cdaql1 ~]$ go_analysis pro
cdaq@cdaql1:~/polhe3/production/hallc_replay$ d2n_status_table 
d2n Collected Data Table [click me to expand]
  • Thu 17 Sep 2020 04:04:21 PM EDT
        Setting      target       P     th  pol  n_runs  IHWP IN      OUT     e-      yield    charge / goal    status 

    Kin-A (hms)   POL-HE3    -4.200  13.50   90     102   51.61%   48.39%  8.57e+07  1.62e+07    5.29 / 6.06     87.3% 
    Kin-A (hms)   POL-HE3    -4.200  13.50  270      17   43.64%   56.36%  1.49e+07  1.36e+07    1.10 / 6.06     18.1% 
    Kin-A (hms)   POL-HE3    -4.200  13.50    0       0                                          0.00 / 0.69      0.0% 
    Kin-A (hms)   POL-HE3    -4.200  13.50  180      23   86.54%   13.46%  1.69e+07  1.42e+07    1.19 / 0.69    170.9% 
    Kin-A (hms)    REF-N2    -4.200  13.50  N/A      14    0.00%  100.00%  1.74e+07  6.99e+07    0.25 / 0.22    113.1% 
    Kin-A (hms)    REF-H2    -4.200  13.50  N/A       8    0.00%  100.00%  3.63e+06  1.24e+07    0.29 / 0.22    133.1% 
    Kin-A (hms)   REF-HE3    -4.200  13.50  N/A      15    0.00%  100.00%  5.57e+06  1.25e+07    0.44 / 0.22    202.1% 
    Kin-A (hms)   REF-VAC    -4.200  13.50  N/A       9    0.00%  100.00%  9.04e+04  4.34e+05    0.21 / 0.22     94.5% 
    Kin-A (hms)   C-OPTICS   -4.200  13.50  N/A      10    0.00%  100.00%  6.40e+06  3.49e+07    0.18 / 0.08    229.3% 
    Kin-A (hms)    EMPTY     -4.200  13.50  N/A       7    0.00%  100.00%  1.33e+06  7.58e+06    0.18 / 0.08    219.6% 
    Kin-B (hms)   POL-HE3    -4.200  16.40   90       0                                          0.00 / 6.13      0.0% 
    Kin-B (hms)   POL-HE3    -4.200  16.40  270       0                                          0.00 / 6.13      0.0% 
    Kin-B (hms)   POL-HE3    -4.200  16.40    0       0                                          0.00 / 0.61      0.0% 
    Kin-B (hms)   POL-HE3    -4.200  16.40  180       8   32.99%   67.01%  1.70e+06  3.94e+06    0.43 / 0.61     70.0% 
    Kin-B (hms)    REF-N2    -4.200  16.40  N/A       6    0.00%  100.00%  2.02e+06  2.89e+07    0.07 / 0.22     31.8% 
    Kin-B (hms)    REF-H2    -4.200  16.40  N/A       0                                          0.00 / 0.22      0.0% 
    Kin-B (hms)   REF-HE3    -4.200  16.40  N/A       6    0.00%  100.00%  9.07e+05  3.42e+06    0.27 / 0.22    120.6% 
    Kin-B (hms)   REF-VAC    -4.200  16.40  N/A       5    0.00%  100.00%  2.74e+04  7.26e+04    0.38 / 0.22    171.8% 
    Kin-B (hms)   C-OPTICS   -4.200  16.40  N/A       2    0.00%  100.00%  4.21e+05  3.53e+06    0.12 / 0.08    149.0% 
    Kin-B (hms)    EMPTY     -4.200  16.40  N/A       2    0.00%  100.00%  1.47e+05  1.14e+06    0.13 / 0.08    161.1% 
    Kin-C (hms)   POL-HE3    -4.000  20.00   90     109   49.29%   50.71%  9.40e+06  1.46e+06    6.45 / 6.05    106.5% 
    Kin-C (hms)   POL-HE3    -4.000  20.00  270     112   53.84%   46.16%  1.12e+07  1.95e+06    5.74 / 6.05     94.8% 
    Kin-C (hms)   POL-HE3    -4.000  20.00    0       0                                          0.00 / 0.69      0.0% 
    Kin-C (hms)   POL-HE3    -4.000  20.00  180      17   69.53%   30.47%  8.33e+05  1.12e+06    0.74 / 0.69    106.8% 
    Kin-C (hms)    REF-N2    -4.000  20.00  N/A      13    0.00%  100.00%  5.04e+06  1.30e+07    0.39 / 0.22    176.8% 
    Kin-C (hms)    REF-H2    -4.000  20.00  N/A       9    0.00%  100.00%  2.58e+05  7.79e+05    0.33 / 0.22    150.7% 
    Kin-C (hms)   REF-HE3    -4.000  20.00  N/A      10    0.00%  100.00%  4.18e+05  9.69e+05    0.43 / 0.22    196.2% 
    Kin-C (hms)   REF-VAC    -4.000  20.00  N/A       6    0.00%  100.00%  3.02e+03  1.48e+04    0.20 / 0.22     92.9% 
    Kin-C (hms)   C-OPTICS   -4.000  20.00  N/A       8    0.00%  100.00%  3.99e+05  7.30e+05    0.55 / 0.50    109.4% 
    Kin-C (hms)    EMPTY     -4.000  20.00  N/A      13    0.00%  100.00%  1.20e+05  5.09e+05    0.24 / 0.50     47.0% 


        Setting      target       P     th  pol  n_runs  IHWP IN      OUT     e-      yield    charge / goal    status 

   Kin-X (shms)   POL-HE3    -7.500  11.00   90     115   48.65%   51.35%  1.47e+08  2.49e+07    5.91 / 6.06     97.5% 
   Kin-X (shms)   POL-HE3    -7.500  11.00  270      72   48.93%   51.07%  7.43e+07  2.39e+07    3.10 / 6.06     51.2% 
   Kin-X (shms)   POL-HE3    -7.500  11.00    0       0                                          0.00 / 0.69      0.0% 
   Kin-X (shms)   POL-HE3    -7.500  11.00  180      27   79.95%   20.05%  3.41e+07  2.37e+07    1.44 / 0.69    207.0% 
   Kin-X (shms)    REF-N2    -7.500  11.00  N/A      25    0.00%  100.00%  4.21e+07  1.28e+08    0.33 / 0.22    150.0% 
   Kin-X (shms)    REF-H2    -7.500  11.00  N/A      11    0.00%  100.00%  6.23e+06  1.82e+07    0.34 / 0.22    155.3% 
   Kin-X (shms)   REF-HE3    -7.500  11.00  N/A      12    0.00%  100.00%  7.50e+06  2.21e+07    0.34 / 0.22    154.6% 
   Kin-X (shms)   REF-VAC    -7.500  11.00  N/A       8    0.00%  100.00%  8.49e+05  3.59e+06    0.24 / 0.22    107.3% 
   Kin-X (shms)   C-OPTICS   -7.500  11.00  N/A      10    0.00%  100.00%  6.40e+06  3.28e+07    0.20 / 0.08    243.8% 
   Kin-X (shms)    EMPTY     -7.500  11.00  N/A      11    0.00%  100.00%  2.66e+06  1.27e+07    0.21 / 0.08    261.9% 
Kin-X-pos (shms)   REF-HE3     7.500  11.00  N/A       3    0.00%  100.00%  7.47e+03  8.19e+04    0.09 / 0.22     41.5% 
   Kin-Y (shms)   POL-HE3    -6.400  14.50   90       0                                          0.00 / 6.13      0.0% 
   Kin-Y (shms)   POL-HE3    -6.400  14.50  270       0                                          0.00 / 6.13      0.0% 
   Kin-Y (shms)   POL-HE3    -6.400  14.50    0       0                                          0.00 / 0.61      0.0% 
   Kin-Y (shms)   POL-HE3    -6.400  14.50  180       7   33.74%   66.26%  2.02e+06  4.91e+06    0.41 / 0.61     66.9% 
   Kin-Y (shms)    REF-N2    -6.400  14.50  N/A       6    0.00%  100.00%  2.16e+06  3.21e+07    0.07 / 0.22     30.6% 
   Kin-Y (shms)    REF-H2    -6.400  14.50  N/A       0                                          0.00 / 0.22      0.0% 
   Kin-Y (shms)   REF-HE3    -6.400  14.50  N/A       6    0.00%  100.00%  1.11e+06  4.32e+06    0.26 / 0.22    117.2% 
   Kin-Y (shms)   REF-VAC    -6.400  14.50  N/A       5    0.00%  100.00%  1.83e+05  4.97e+05    0.37 / 0.22    167.4% 
   Kin-Y (shms)   C-OPTICS   -6.400  14.50  N/A       2    0.00%  100.00%  4.35e+05  3.67e+06    0.12 / 0.08    147.9% 
   Kin-Y (shms)    EMPTY     -6.400  14.50  N/A       2    0.00%  100.00%  2.02e+05  1.57e+06    0.13 / 0.08    160.4% 
   Kin-Z (shms)   POL-HE3    -5.600  18.00   90      99   52.63%   47.37%  6.77e+06  1.20e+06    5.62 / 6.05     92.9% 
   Kin-Z (shms)   POL-HE3    -5.600  18.00  270      68   52.79%   47.21%  4.58e+06  1.18e+06    3.88 / 6.05     64.0% 
   Kin-Z (shms)   POL-HE3    -5.600  18.00    0       0                                          0.00 / 0.69      0.0% 
   Kin-Z (shms)   POL-HE3    -5.600  18.00  180      19   60.72%   39.28%  9.41e+05  1.08e+06    0.87 / 0.69    125.2% 
   Kin-Z (shms)    REF-N2    -5.600  18.00  N/A       3    0.00%  100.00%  1.19e+06  5.66e+06    0.21 / 0.22     95.8% 
   Kin-Z (shms)    REF-H2    -5.600  18.00  N/A      10    0.00%  100.00%  2.65e+05  8.22e+05    0.32 / 0.22    146.7% 
   Kin-Z (shms)   REF-HE3    -5.600  18.00  N/A      10    0.00%  100.00%  4.24e+05  1.01e+06    0.42 / 0.22    190.7% 
   Kin-Z (shms)   REF-VAC    -5.600  18.00  N/A       6    0.00%  100.00%  1.75e+04  8.85e+04    0.20 / 0.22     90.0% 
   Kin-Z (shms)   C-OPTICS   -5.600  18.00  N/A       7    0.00%  100.00%  2.74e+05  5.19e+05    0.53 / 0.50    105.7% 
   Kin-Z (shms)    EMPTY     -5.600  18.00  N/A       7    0.00%  100.00%  3.90e+04  2.71e+05    0.14 / 0.50     28.8% 
Kin-Z-pos (shms)   REF-HE3     5.600  18.00  N/A       0                                          0.00 / 0.22      0.0% 


  • The above table is also posted to the HCLog every shift with the subject "Recent SHMS and HMS run list".
    • You can use the Hall C Run List Summaries link under the HCLog Useful Links sidebar on the right as a shortcut.
  • The runtimes and estimated yields are based on Rate Estimate Updates (19 Mar 2020) by Wolfgang
    • (NOTE: Ignore the primed kinematics in the file, they are not relevant anymore.)

Special Runs

Miscellaneous Calibrations

Target Calibrations and Tests

 A) 90 degree
   1) Spin-up in various conditions (Done for some conditions)
   2) EPR calibration (done)
   3) PNMR initial calibration (done)
      3a) Follow up PNMR/NMR calibrations (done during NMR measurements)
   4) AFP losses for various conditions (done for one condition)
   5) Temperature test (done)
   6) Optimization of QWPs (done)
   7) Studies/Optimizations of lasers/temperature/convection speed (partially done)
   8) Cross calibrations of NMR at different target positions (mainly pickup and He3-in-beam) (optional, can wait to after run ends)
   9) Cross calibrations for different HB settings (3 settings for 5 pass) (optional, can wait to after run ends)
   10) Beam effects study (from regular polarimetry data?)
 B) 270 degree
   1) Spin-up in various conditions
   2) EPR calibration (done)
   3) PNMR initial calibration (done)
      3a) Follow up PNMR/NMR calibrations
   4) AFP losses for various conditions
   5) Temperature test
   6) Optimization of QWPs
   7) Studies/Optimizations of lasers/temperature/convection speed
   8) Cross calibrations of NMR at different target positions (mainly pickup and He3-in-beam)
   9) Cross calibrations for different HB settings (3 settings for 5 pass)
   10) Beam effects study
   11) Cross calibrations between 90 and 270

Optics Running (OPTIONAL)

  • We took all of the critical Optics data we need back in March. (These have been analyzed.)
  • [OPTIONAL] Additional Optics running
    • Kin-X (SHMS) P = -5.600, Th = 18.00 (Sieve, No-Sieve) (6--8 hours)
    • Kin-B (HMS) P = -4.200, Th = 16.40 (No-Sieve) (1 hour)
    • NOTE: (We would take the above runs together and likely just take both Sieve

and No-Sieve in each arm.

Notes for Near Term Running

NOTE: This will be updated frequently

Positron Running (Charge Sym BG study)

F2 CSB Data LH2 (or LD2?)
Marathon CSB Data on He3

- Simona showed a nice plot from the Hall C F2 running of their CSB correction at 21deg vs E'.

 That indicates the e+:e- correction is indeed quite small (she'll send out the F2 data to the group)
 - That argues that all we need to measure are rates -- much simpler/faster.

- Mike Nycz provided some data from Marathon as well

- For rates only, it looks like it would be better to use the SHMS instead since its polarity can be reversed remotely

 - We will juggle the positive running in Kin-Z with taking longitudinal data on Thu/Fri (Aug 20, 21)
 - Warning (Aug 20):  The positive SHMS running seems to have impacted the target much more than anticipated.
 - Took positive SHMS running for Kin-X (+7.500  11.00deg) on OWL, Aug 20 -- Done
 - WANT positive SHMS running for Kin-Z (+5.600  18.00deg) -- TBD

Longitudinal Running

 - Every kinematic requires roughly 10--12 hours of longitudinal running
 - We will attempt some longitudinal running for this kinematic pair (Kin-A, Kin-X) starting
     Thursday swing following the Moller (Aug 20)
 - We will have to come back to these kinematics next week!

Reference Cell Running

 - Every kinematic requires roughly 12 hours of Reference cell running
   - REF-Hydrogen, REF-Nitrogen, REF-Vacuum, REF-He3
   - NOTE: Make He3 the last ref-cell fill before a major change, so we 're-use' the He3
           in the next setting.

Pol 3-He Cell Changes

See also: Production Target Cell Information

  • Replaced Austin with Briana on Aug 21--24
  • Tommy being remounted; will be next available cell
  • Butterball will be prepped next
    • NOTE: Butterball is quite short (21mm short vs. 3mm tolerance; associated issues must be addressed)
Plan is to install a new cell following the 1-pass running (start Thurs, Aug 3?)

1-pass Running • Starting Fri Sept 18; Ending Mon Sept 21

 - See D2n_Run_Plan_2020#1-pass_Running below for details

5-pass Production Procedure

  • SHMS and HMS runs are independent, but are run in pairs that should reflect similar rates (when possible)
    • Run time for each of the 3 kinematic pairs is roughly 125 PAC hours, ~10 calendar days/pair
  • Minimize the number of NMR measurements to preserve target polarization
    • This generally means minimizing target direction changes, since each change usually imposes an 'as-is' NMR, and then a post-rotation NMR.
  • General considerations
- Flip IHWP every 4--8 hours (coordinate with Hall A)
- Ensure Charge Asym Feedback is running/working (watch the stripcharts and coordinate with Hall A)
- Take NMR meas. every ~5 hours
- Schedule a Moller run roughly every week
  - Target spin rotations/tests should be done immediately prior to the Moller so target has time to spin back up
  • Reference cell runs needed for all Kinematics (nominal 2--4 hours per kinematic)
 - Reference cell Nitrogen [135 psig]  (Insert when we need time to spin up Polarized Cell)
 - Reference cell Hydrogen [135 psig]  (Insert when we need time to spin up Polarized Cell)
 - Reference cell Vacuum   [0   Torr]  (Insert when we need time to spin up Polarized Cell)
 - Reference cell He3      [135 psig]  (NOTE: Make best use of He3 by running this last in first kinematic, then first in next kinematic!)
  • Polarized 3He Target • Transverse (90, 270) • Total run time 8--10 days
- Flip target spin after 3--4 days
- Run 3--4 more days with the opposite target spin
  • Polarized 3He Target • Longitudinal (0, 180) • Total run time ~1--2 days
 - Flip IHWP at 1/2 way into each run block (~4 hours if possible)
  - ~11 hours @   0deg
  - ~11 hours @ 180deg

1-pass Running

         Still cleaning this up, but all the major items should be represented here.
  • Pass change to 1-pass starts at 6am, Fri Sept 18th, and continues to end of program at 8am on Monday, Sept 21
 NOTE Run Plan for Shift Crew: D2n_Run_Plan_2020-1pass 

Priorities for 1-pass running

Rough priority list for 1-pass running. 4, 2 are most important. Unpolarized running beam energy measurement are front-loaded while cell polarizes.

  1. Beam energy measurement
  2. Cross section meas. on reference cell
    1. Elastic H2, He3 (nominal 60 min each) (2--3 hours)
    2. Cross section for empty, Carbon (30--60min each; 2 hours)
  3. DIS (flat spectrum) acceptance check on H2 (1--2 hours)
  4. Longitudinal 3He elastic asym (target 3% stat err after cuts) (8--16 hours)?
    1. Chao estimates 16hrs : dA/A of 2.3%; 8hrs : 3.2% with the paddles set as planned. We'll take a couple runs with all-paddles on to verify paddle settings and to get a full acceptance sample.

  1. pressure curves for N2 and He3
  2. Transverse delta asym

  1. Elastic 3He program

Verify the following before 1-pass period

  • Collimator is located and ready to go; installation plan is clear;
  • Determine if EPR photo-diode will need replacement -- NO, we will not (Junhao checked and it should still be OK if we need it. Can always replace post run too.)
  • Starting angles: SHMS: 8.5 deg; HMS 11.7 deg
  • Get the 1-pass correction coil settings from Jixie -- Done
  • Confirm the charge-asym feedback system will be up and running while Hall A is off -- Done
  • Ensure the 1-pass analysis scripts are ready (Scott, Chao); Chao reports that is ready to go now
  • All set to switch Raster to 1-pass mode (Gunning contacted, MCC knows, have instructions from Mark)
  • Update run-time estimates for long. 3He elastics at 8.5 deg (Chao)
    • Chao estimates 16hrs : dA/A of 2.3%; 8hrs : 3.2% with the paddles set as planned. We'll take a couple runs with all-paddles on to verify paddle settings and to get a full acceptance sample.
  • Review the 'Scint. bar configuration' notes below
Notes on 1-pass elastic scripts [click me to expand]

[Notes from Chao] Subject: RE: Online Plots for 1-Pass Elastic Experiment

For performance I have temporarily commented the elastic analysis in the scripts:

in "hallc_online/analysis/hcfull":

   script_file = '../hallc_online/analysis/elastic/el_asym_{}.cxx'.format(args.command.lower())
   script_args='{},{},\"{}\"'.format(args.run, -1, args.mode)
   analysis_cmd = "root -b -q '{}+({})'".format(script_file, script_args)
  1. os.system(analysis_cmd) <- this line

in "hallc_online/util/post_plots_to_logbook": -a "${ODIR}"/MON_OUTPUT/ANLYSIS_HISTOS/${2}/PID_variables2.png

   # -a "${ODIR}"/MON_OUTPUT/ANLYSIS_HISTOS/${2}/elastic1.png       <- this line
   # -a "${ODIR}"/MON_OUTPUT/ANLYSIS_HISTOS/${2}/elastic2.png       <- this line

Please uncomment them for the auto-posting

The elastic asymmetry runs were determined with the momentum setting (-p) and angle (-a) setting in "hallc_online/bin/make_elastic_asym_plots":

python3 analysis/elastic/asym_elastic.py HMS -p 2.17 -a 11.7 -t 0.1 --estimate=0.0503 --factor-qe=0.65 --date-start=2020-08-01

python3 analysis/elastic/asym_elastic.py SHMS -p 2.17 -a 8.5 -t 0.1 --estimate=0.026 --factor-qe=0.80 --date-start=2020-08-01

here "--date-start" is used to separate d2n elastic runs from a1n elastic runs.


The scripts for elastic runs are ready, and I have added them for the auto-posting.

There will be two elastic plots in the analysis post for each run, showing the elastic cuts.

These posted plots are generated by the scripts

"~/polhe3/production/hallc_online/analysis/elastic/el_asym_(s)hms.cxx"

At the end of each shift, the summary post (recent SHMS and HMS run list) will show the total asymmetry from all the elastic runs.

Currently I am using the A1n elastic settings, please let me know if any changes are needed.

Or you can change the settings in "~/polhe3/production/hallc_online/bin/make_elastic_asym_plots", as:

python3 analysis/elastic/asym_elastic.py HMS -p 2.17 -a 11.7 -t 0.1 --estimate=0.0503 --factor-qe=0.65 --date-start=2020-08-01

python3 analysis/elastic/asym_elastic.py SHMS -p 2.17 -a 8.5 -t 0.1 --estimate=0.026 --factor-qe=0.80 --date-start=2020-08-01

Please check the following instruction for the script's arguments:

positional arguments: arm Spectrometer Arm (HMS or SHMS)

optional arguments:

 -h, --help           show this help message and exit
 -p MOMENTUM          momentum setting (GeV)
 -a ANGLE             angle setting (degree)
 -t TOLERANCE         absolute tolerance for settings
 --date-start START   start date
 --date-end END       end date
 --counts-min CMIN    minimum event counts for valid elastic runs
 --estimate ESTIMATE  estimated asymmetry for this setting (without any corrections)
 --factor-qe F_QE     dilution factor from Quasi-Elastic asymmetries
 --run-script         run count script if output does not exist

Necessary Prep work during Hall Controlled Access on Monday (schedule with RadCon in advance)

  • Install collimators (JP, RadCon) [2]
  • Check on EPR photo-diode status, likely replace it (RadCon)
  • Let Hall B know that we will be changing IHWP status more often during this period
    • Likely every 2--4 hours (not 8 hours)
    • We will still let them know in advance.
  • Ensure fast raster is reconfigured correctly (Expert: Jones, Gunning)
    • "Have MCC click on the button on the Expert page for 1 coil (instead of 2 coils). Ensure "gbeam_pass1_jan2020.param" is used for analysis"
    • Bill G support will support shut off / unplug one of the power supplies (Brad will follow-up)
    • Verify pattern using DAQ and random pulser (Brad)
  • Rotate spectrometers to minimum angles for HMS (11.7) and SHMS (8.5) (S. Lassiter, A. Comer)
    • Confirm we can rotate them away from the beamline remotely
  • Pressure curve questions
    • Pump Reference cell to vacuum while waiting for Sweep in the morning
    • Open He3 bottle just before locking up Hall
    • Complete reference cell runs after we come back up (following Beam energy meas.)
    • Escorted access to Hall after ref cell running complete to close valves at He3 cylinder (Opportunistic, or should we stop the program to do this?)
  • Confirm SHMS/HMS momentum settings
  • Set target to Longitudinal/180
    • Do we want to attempt flip to Longitudinal/0 during 3He elastics?

SHMS/HMS scintillator bars settings for He3 Elastic and QE runs

SHMS Paddle Configuration Optimized for He3 Elastic Measurement (8.5deg, -2.12860 GeV/c)

(From first_epics_HV.results https://logbooks.jlab.org/entry/3762329)

  • Turn OFF all 1x paddles ((both sides: '+' and '-') EXCEPT 07, 08
  • Turn OFF all 2x paddles ((both sides: '+' and '-') EXCEPT 07, 08, 09
    • Only this set of s1x and s2x channels should be ON
   s1x07+(L) s1x07-(R)
   s1x08+(L) s1x08-(R)
   
   s2x07+(L)  s2x07-(R)
   s2x08+(L)  s2x08-(R)
   s2x09+(L)  s2x09-(R)
  • Keep all y paddles ON (unchanged)

HMS Paddle Optimization for He3 QE Measurement

  • Turn OFF paddles 01 through 06 in h1x (both sides: '+' and '-')
  • Turn OFF paddles 01 through 06 in h2x (both sides: '+' and '-')
  • Only this set of HMS h1x and h2x paddles should be ON
    h1x07(+) through h1x16(+)
    h1x07(-) through h1x16(-)
    
    h2x07(+) through h2x16(+)
    h2x07(-) through h2x16(-)
  • Keep all y paddles ON (unchanged)

HMS Paddle Config Optimized for He3 Elastic Measurement (11.7deg, -2.08234 GeV/c)

NOTE: These are a little more restrictive (more bars off); probably want to start with the above looser setting and then fine-tune (From first_epics_HV.results https://logbooks.jlab.org/entry/3762316)

  • Turn OFF paddles 01 through 09 in h1x (both sides: '+' and '-')
  • Turn OFF paddles 01 through 09 in h2x (both sides: '+' and '-')
  • Only this set of HMS h1x and h2x paddles should be ON
    h1x10(+) through h1x16(+)
    h1x10(-) through h1x16(-)
    
    h2x10(+) through h2x16(+)
    h2x10(-) through h2x16(-)

A1n paddle studies for 3He elastics in SHMS, and QE in HMS

  • For reference

Run Plan for 1-pass

Beam checkout

Moller Measurement (4--8 hours; Cancelled)

Consensus is to skip the Moller at 1-pass.
  • Notes from Dave Gaskell
If we don't change the Wien angle for 1 pass running, Hall C should
receive 97.5% of maximum polarization (which is about 87.5% now, so Hall
C would get 85.3%).

At 1 pass and close to maximum polarization, we are very insensitive to
absolute beam energy and Wien angle.  Even incorporating a 3 degree
uncertainty in the Wien angle would only change the polarization about
1.3%. Few MeV differences in the beam energy, much less.

Anyway - I don't think there's much risk in skipping the 1 pass Moller
measurement in Hall C.  A super-conservative estimate of the
polarization uncertainty would be 4%. Conservative: 3%.
  • Beam parameters to check at 1-pass:
    Photocathode QE         0.45280     (qe_hallc; nominal, should be above 0.30)
    Wien Angle:horizontal  -29.6402     (HWienAngle;  should NOT change)
    Beam energy             2179.7 MeV  (HALLC:p; nominal value is 2183 MeV)

Beam Energy Measurement (1 hour)

  • Accelerator driven process -- will happen following initial beam checkout into the Hall
  • Would be recommended, since it is a sensitive input to the beam polarization that Hall C receives

Reference Cell Hydrogen (H2) Cross Section (1--2 hours)

 Goals 
 - To measure hydrogen elastic cross section as a cross check of data quality
 - To measure backgrounds from glass cell and beamline windows (Reference-VAC)
 Setup
 - Beam Current: up to 30uA and DAQ rate-limited, raster always ON (4mm diameter)
 - Target:       Reference Cell: vacuum and filled with H2 to 135 psig
 - Detectors:    All scintillator bars on
 Spectrometer Setup (with HMS 11.7 deg, SHMS 8.5 deg)
 - HMS setting: 11.7 deg, -2.08234 GeV/c (electron QE)  -- Jan 3He elastic settings
 - SHMS setting: 8.5 deg, -2.12860 GeV/c (electron QE setting, but focus on 3He elastic events) -- Jan 3He elastic settings
 Procedure
 - Set spectrometers at the required settings, target at the Reference Cell position;
 - Evacuate Reference Cell (to vacuum), record the pressure reading in the Start_Run entries.
 - Ask MCC for 10uA raster ON (4mm dia), check rates and prescale factors
   (Save PS file if changed). Goal is to reach (3.5-4)kHz DAQ rate on both HMS
   and SHMS. If rates too low, increase the beam current but do not exceed
   30uA. If rates too high, prescale it down and do NOT drop current below
   10uA.
 - Take one 20-minute run, make sure there are no less than 10,000 events for each of HMS and SHMS.
 - Record the pressure in End_Run entries. 
 - Fill Reference Cell to 135 psig of hydrogen gas;
 - Keep beam current the same as the above Empty Cell run, check rates and
   prescale factors (Save PS file if changed). Goal is to reach (3.5-4)kHz DAQ
   rate on both HMS and SHMS. If rates too low, increase the beam current but
   do not exceed 30uA. If rates too high, prescale it down and do NOT drop
   current below 10uA.
 - Take one 20-minute run, apply ztar and W cut and make sure H2 elastic
   events from the H2 gas are no less than 10,000 for each of HMS and SHMS.
 - Record the pressure reading in the Start_Run and End_Run entries.
 Things to watch online
 - Ensure all scintillator bars are ON
 - For 10atm H2 filled cell, apply ytar or ztar cuts to isolate events
   scattered from H2 gas (We do not want events from the glass windows!), and
   apply W cuts to select only H2 elastic. Note the elastic peak is not at
   dp=0 because the spectrometers are set at 3He elastic. 
 - Run online script to ensure we collect no less than 10,000 H2 elastic (with
   ytar or ztar and W cuts) for each of HMS and SHMS

3He Reference Cell Runs (N2 and 3He Pressure Curves, False Asymmetries) (5hrs ideal, 3 hrs min)

 Goals 
 - For empty cell, to determine window background for elastic setting
 - For N2, need 5 points at 0.5(*), 1.0, 2, 4 and 8 atm, 10,000 N2 elastic events for each pressure setting and each spectrometer.
   (*) sub-atmosphere may not be possible (gauge not accurate enough)
 - For 3He, need 6 points at 1, 2, 4, 6, 8 and 10 atm, 10,000 3He elastic events for each pressure setting and each spectrometer.
 - Reference cell must be filled with 3He last, and do NOT pump the 3He gas out!
 - Time: couple of hours assuming 50% beam efficiency
 Setup
 - Beam Current: 10-30uA, raster always ON (4mm diameter). Do not drop currents below 10uA.
 - Target:       Reference Cell, Empty, filled with N2, filled with 3He (last)
 - Detectors:    We will need to keep only 2 scintillator bars ON to take mostly 3He elastic and limit quasi-elastic scattering events. 
 - Use these settings: Scintillator Settings
   A1n notes: hclog 3756736, hclog 3760238
   - HMS Detectors: turn off paddles 1-6 on 1X plane and paddles 1-6 on 2X plane
   - SHMS Detectors: keep only paddles 7,8 of 1X plane, and paddles 7,8,9 of 2X ON, all others off;
   - SHMS (continued): Also suggest for 1Y keep only 7 on, all others off, although this will depends on how well the collimators works


 Spectrometer Setup (with HMS 11.7 deg, SHMS 8.5 deg)
 - HMS setting: 11.7 deg, -2.08234 GeV/c (electron QE)  -- Jan 3He elastic settings
 - SHMS setting: 8.5 deg, -2.12860 GeV/c (electron QE setting, but focus on 3He elastic events) -- Jan 3He elastic settings
 Procedure
 - Set spectrometers at the required settings, target at the Reference Cell position.
 - Make sure the correct scintillator bars are ON.
 - Start at for 10 uA raster ON (4mm dia), check trigger rates and PS factors.
   Ideally we would like to see (3.5-4)kHz DAQ rates. If rates are too high,
   use Prescaling (save PS file if changed). If rates are too low, increase
   beam current but do not exceed 30uA.  Ensure that the livetimes are >80%
   in each arm.
 Nitrogen Pressure Curve
 - Evacuate cell to vacuum and take one 20-minute run.
 - Fill the cell (slowly) to 380 Torr with N2, take one 20-minute run.  (NOTE: vacuum gauge is unreliable due to overpressure exposure.  Is this useful?)
 - Fill the cell (slowly) to 1 psig, take one 20-minute run.
 - Fill the cell to 15 psig, take one 20-minute run.
 - Fill the cell to 45 psig, take one 20-minute run.
 - Fill the cell to 75 psig, take one 20-minute run.
 - Fill the cell to 105 psig, take two 45-min runs and flip the IHWP in between
 He3 Pressure Curve
 - Evacuate cell to vacuum and take one 20-minute run.
 - Fill the cell (slowly) to 380 Torr with He3, take one 20-minute run. (NOTE: vacuum gauge is unreliable due to overpressure exposure.  Is this useful?)
 - Fill the cell (slowly) to 1 psig, take one 20-minute run.
 - Fill the cell to 15 psig, take one 20-minute run.
 - Fill the cell to 45 psig, take one 20-minute run.
 - Fill the cell to 75 psig, take one 20-minute run.
 - Fill the cell to 105 psig, take one 20-minute run.
 - Fill the cell to 135 psig, take two 45-min runs and flip the IHWP in between
 - Important: After these runs are completed, do NOT pump out the 3He gas!
 Things to watch for
 - These reference cell runs must be taken with exactly the same spectrometer and scintillator settings as polarized 3He elastic asymmetry runs (see next).
 - Beam current should not be below 10uA (BCM accuracy) and should not be above 30uA (glass cell safety).
 - For each gas pressure setting (except for empty cells), need 10,000 events from gas alone (apply ztar cuts) for each of HMS and SHMS.

Polarized 3He Cell Density Measurement (1 hour plus time for setup)

 Goals 
 - To determine 3He and N2 density in cell Tommy
 
 Setup
 - Beam Current: 10uA, raster always ON (4mm diameter)
 - Target:      Polarized 3He Cell "Tommy", polarized in 180 deg.
 Spectrometer Setup
 - HMS setting: 11.7 deg, -2.1483 GeV/c (electron 3He elastic) for density of "Tommy".
 - SHMS setting: 8.5 deg, -2.1286 GeV/c (electron QE) for collimator check and asymmetry. 
 - SHMS Detectors: follow QE/el measurement setup below.
   -  HMS Detectors: turn off paddles 1-6 on 1X plane and paddles 1-6 on 2X plane
   - SHMS Detectors: keep only paddles 7,8 of 1X plane, and paddles 7,8,9 of 2X ON, all others off;
   - SHMS (continued): Also suggest for 1Y keep only 7 on, all others off, although this will depends on how well the collimators works
   - See Scintillator Settings
 Procedure
 - Set spectrometers at the required settings, target at the Polarized 3He Cell position;
 - Check that scintillator bars in HMS and SHMS are as specified above.
 - Ask MCC for 10 uA raster ON (4mm dia). Adjust prescale factor to limit the DAQ rate to 3.5-4kHz. 
   (Do NOT drop below 10uA. If rates are low, increase beam current but do not exceed 30 uA.)
 - Take two 30-min runs at this setting. (see below)
 Things to watch for
 - We need at least 10k events each from 3He elastic and N2 elastic peaks.
 - Run Chao's script to pull out # of events for 3He elastic and N2 elastic and post in hclog;
 - If we don't expect to achieve 10k events within one hour, contact RC.
 - For SHMS, post ytar(ztar) histogram in hclog. This shows effects of the new collimator.
 - Chao will run analysis remotely to extract 3He elastic fraction on SHMS.

Polarized 3He Runs, 3He elastic and QE longitudinal Asymmetries (12--16 hours)

 Goals 
 - To measure physics longitudinal asymmetry of 3He elastic (on SHMS) and QE (on HMS) scattering. 
 - If precision is high, this cross checks PbPt from Moller and Target Polarimetry. 
 - If precision is low, at the minimum we should determine the sign of asymmetries.
 - Experts must watch EDTM rate (Brad).
 Setup
 - Beam Current: 5-30uA, raster always ON (4mm diameter)
 - Target:      Polarized 3He Cell
 - Target spin direction: 180 deg
 - See Target Field Runplan for coil currents settings.
 - Time: total 12--16 hours of data taking, flip beam IHWP every 4 hours. Do
   target NMR every 4 hours (ask TO), but always before and after the full data taking.
 Spectrometer Setup (with HMS 11.7 deg, SHMS 8.5 deg)
 - HMS setting: 11.7 deg, -2.08234 GeV/c (electron QE)  -- January setting
 - SHMS setting: 8.5 deg, -2.12860 GeV/c (electron QE setting, but focus on 3He elastic events) -- January setting
 - Use these settings: Scintillator Settings
   A1n notes: hclog 3756736, hclog 3760238
   - HMS Detectors: turn off paddles 1-6 (or 1-9) on 1X plane and paddles 1-6 (or 1-9) on 2X plane
   - SHMS Detectors: keep only paddles 7,8 of 1X plane, and paddles 7,8,9 of 2X ON, all others off;
   - SHMS (continued): Also suggest for 1Y keep only 7 on, all others off, although this will depends on how well the collimators works
 Procedure
 - Set target at the Polarized 3He Cell position, spin 180 deg.
 - Perform one full set of polarimetry (NMR, EPR if possible) measurements before taking data.
 - Ask MCC for 10 uA raster ON (4mm dia), check trigger rates. Goal is to reach (3.5-4)kHz DAQ rates.
 - If rates too high, adjust PS factors (save PS file if changed!) and/or drop beam current but do not go below 5uA.
 - If rates too low (unlikely), increase beam current but do not exceed 30uA. 
 - Perform NMR every 4--5 hours
 - Start with all bars ON
 - Target is set to Longitudinal/180
 - Take 1 60-minute run (assuming 50% beam efficiency) (ALL BARS ON)
   - 30 minutes in, start analysis and verify above paddle settings are optimal
 - Configure HMS scintillator bar status the same as suggested above
 - Configure SHMS scintillator bar status the same as suggested above
 - Take 3 60-minute run (assuming 50% beam efficiency)
 
 - Call MCC and ask beam IHWP to be inserted (or taken out). Need to inform other halls prior to the switch.
 - Take 4 60-minute runs (assuming 50% beam efficiency).
 - Call MCC and ask beam IHWP to be inserted (or taken out). Need to inform other halls prior to the switch.
 - Take 2 60-minute runs (assuming 50% beam efficiency).
  
 - Rotate target to Longitudinal/0
 - Turn all bars ON
 - Take 1 60-minute run (assuming 50% beam efficiency) (ALL BARS ON)
 - Configure HMS scintillator bar status the same as suggested above
 - Configure SHMS scintillator bar status the same as suggested above
 - Take 3 60-minute run (assuming 50% beam efficiency)
 Things to watch online
 - Ensure the correct scintillator bars are on (We need 3He elastic events in SHMS and QE events in HMS)
 - Apply z cuts to isolate events scattered from 3He gas (We do not want events from the glass windows!)
 - Apply W or dp cuts to isolate 3He elastic (SHMS) or QE(HMS) events from others.
 - For HMS, run online script to ensure we see physics asymmetry. Plot asymmetry as function of nu (=E-E'). 
 - For SHMS, run online script to ensure we see physics asymmetry. Plot asymmetry as function of W(3He). 
 Predictions
 - QE asymmetry (HMS) predictions are given in hclog 3759756. 
 - Elastic 3He asymmetry (SHMS) is predicted to be 2.5% on SHMS, with Pb=0.8, Pt=0.5 and fN2=0.9 gives Ameas=0.9%.
 - Important: would like to have more than 6% of the full rate on SHMS to come from 3He elastic.

Polarized 3He Delta Transverse Asymmetry (4--8hr)

 Goals 
 - To determine the sign of transverse asymmetry of Delta resonance on HMS and SHMS.
 - On SHMS we will also take QE events which help to determine sign of transverse asymmetry.
 - Experts must watch EDTM rate (Brad or Sylvester)
 Setup
 - Beam Current: 5-30uA, raster always ON (4mm diameter)
 - Beam IHWP:    should be switched half-way
 - Target:       Polarized 3He Cell
 - Detectors:    All scintillator bars ON.
 - Parity check: Inform Hall A to make sure they set the Qasym feedback cutoff at 2uA.
 - 
 Spectrometer Setup (with HMS 11.7, SHMS 8.5 deg)
 - HMS setting:  11.7 deg, -1.75830 GeV/c (electron W=1.232)
 - SHMS setting:  8.5 deg, -1.79736 GeV/c (electron W=1.232)
 Pol 3He Target Setup (TO and Target Expert)
 - Target spin direction: pointing 270
 - Ensure correction coils are set properly
 Procedure
 - Set target at Polarized 3He Cell position, spin pointing either beam-left or beam-right.
 - Check and make sure all scintillator HVs are ON (both HMS and SHMS).
 - Perform one set of polarimetry (NMR, EPR if possible) measurements before taking data.
 - Ask for 5uA raster ON, check rates and set prescale factors (Save them if changed!). 
 - Goal is (3.5-4)kHz DAQ rates on both HMS, SHMS. 
 - If rates too low (unlikely), can increase current to reach the desired rates but do not exceed 30uA.
 - If we have 4 hours of beam time:
 - Take two 60-minute (beam ON at least 30 min each) runs.
 - Switch beam IHWP position (need to inform other halls in advance). Either from IN to OUT or from OUT to IN.
 - Take two 60-minute (beam ON at least 30 min each) runs.
 - Perform one set of polarimetry (NMR, EPR if possible) measurements.
 - If we have 8 hours of beam time:
 - Take four 60-minute (beam ON at least 30 min each) runs.
 - Perform one NMR measurement. Confirm target polarization
 - Switch beam IHWP position (need to inform other halls in advance). Either from IN to OUT or from OUT to IN.
 - Take four 60-minute (beam ON at least 30 min each) runs.
 - Perform one set of polarimetry (NMR, EPR if possible) measurements.
 Things to watch online
 - Apply y or z cuts to isolate events scattered from 3He gas.
 - Apply 1.1<W<1.35 GeV cut to isolate the Delta resonance.
 - Run online script to ensure we see physics asymmetry. Plot asym vs. W and nu. 
 - HMS and SHMS should see opposite signs of asymmetries (but not the same magnitudes). 
 - Expected asymmetries for Delta(1232): Ameas=0.67% at 11.7deg and 0.46% at 8.5 deg if Pt=50%, Pb=85%, N2 dilution 0.92.
 - Expected asymmetries for QE are given in hclog 3759756 (ignore Aperp2,
   SHMS will see Aperp always, up to a sign)
 - Expected rate: 5kHz at 1uA, about 10% (5% w/o collimators) would be
   Delta(1232) from 3He gas for HMS (SHMS), see hclog 3758839 analysis of 11.7
   deg data taken previously.
 - Counts needed to reach 10% on Ameas relative: 2.2M on HMS or 4.7M on SHMS.
   If Pt below 50%, scale up by (50%/Pt)^2.
 - Expected time to reach 10% on Ameas relative: 1.6 pac hours for HMS, 5.6
   PAC hours for SHMS with all paddles on and w/o collimators.  Adding target
   collimators on the SHMS side will help significantly).


Polarized 3He Elastic Longitudinal Asymmetry (until end of 1-pass running)

3He Asyms for 2.18 GeV/c (S. Barcus, S. Li)
  • Continue with the Pol He3 Elastic group's run plan and continue stepping the SHMS out in angle) until we run out of beam time.
    • Rotate target back to 180deg
    • SHMS set to 11.0 deg to start, -2.08234 GeV/c (? Confirm)
    • HMS will rotate to 18.5 deg, -2.12860 GeV/c (? Confirm) and park there for remaining 3He Elastic Asym running
  • Settings to be confirmed by Shujie, Scott

3He Asym rates and estimates for 2.18 GeV/c (S. Li, S. Barcus)

  • new runplan (0918) TBD
  • Note: These values are all estimates. Optimize current and prescale values (using the el_clean trigger ps3) to produce a data rate of ~4 kHz!
1 pass runplan
Kinematics Angle (degree) Ep (GeV) Current (uA) length (beam on hours) Prescale Setting (PS3) estimated rate (total)
HMS Kin-1A 11.7 2.0823 5 D2n Systematics 1 2219
HMS Kin-1D 17 2.0823 30 Remaining time 0 2003
SHMS Kin-2A 8.5 2.1289 5 D2n Systematics 3 3045
SHMS Kin-2C 11 2.1289 5 2 1 2643
SHMS Kin-2D 15 2.1289 20 2 0 3971
 Goals 
 - To measure physics longitudinal asymmetry of 3He elastic scattering to (2-3)% relative, cross check PbPt from Moller and Target Polarimetry
 Setup
 - Beam Current: 10-30uA, raster always ON (4mm diameter)
 - Target:      Polarized 3He Cell, Reference Cell
 - Target spin direction: 180deg only
 - See Target Field Runplan for coil currents settings.
 - Detectors:    We will need to keep only a few scintillator bars ON to accept 3He elastic and limit quasi-elastic scattering events. 
   - HMS Detectors: turn off paddles 1-6 on 1X plane and paddles 1-6 on 2X plane
   - SHMS Detectors: keep only paddles 7,8 of 1X plane, and paddles 7,8,9 of 2X ON, all others off;
   See Scintillator Settings
 - Parity check: Inform Hall A prior to taking data to make sure they set the charge asymmetry feedback cutoff at 2uA.
 - Time: total 8 hours of data taking with even split between IHWP settings
   Full target polarimetry measurements at the start and the end, plus pNMR
   (pulsed NMR) only every two hours or after each target spin flip.
 Procedure
 - Verify the correct Scintillator bars are ON.
   - Scintillator Bar Settings
 - Move target to Polarized 3He Cell position, spin polarization at 180 deg.
 - Perform one full set of NMR, (EPR?), pNMR polarimetry measurements before taking data, and every 4 hours thereafter
 - Set SHMS and SHMS DAQs to use EL_clean (T3) triggers unless unprescaled T1 rates are < 3.5 kHz
 - Ask MCC for 10uA raster ON (4mm dia), check trigger rates. Goal is to reach
   3.5kHz--4kHz DAQ rates. If rates too high, adjust PS factors (save PS file if
   changed!). If rates too low (unlikely), increase beam current but do not
   exceed 30uA.
 - Want at least 2--4 hours of beam on target for inital(8.5 deg) point; flip IHWP at half-way mark
 Things to watch online
 - Ensure the correct scintillator bars are on (We need 3He elastic events!)
 - Apply z cuts to isolate events scattered from 3He gas (We do not want events from the glass windows!)
 - Apply W or dp cuts to isolate 3He elastic events from others.
 - Run online script to ensure we see physics asymmetry. HMS and SHMS should
   see the same asymmetry. Asymmetry should flip sign with beam IHWP switch
   and target spin flip.
 - Expected asymmetries (3He elastic only): Aphys=5.0% Ameas~2% at 11.7 deg, or Aphys=5.8% Ameas=2.3% at 12.5 deg, if using Pt=50%, Pb=85%, N2 dilution 0.92.
 - Expected uncertainties dA/A=2% (goal) or 3% (minimum).

Important Scripts

  • Elastic Electron Count Script:
    • Location: /home/cdaq/polhe3/production/hallc_online/analysis/elastic/el_asym_(s)hms.cxx
    • Description: These scripts count the number of elastic electrons in a data run after applying various cuts.
    • Usage: To use these scripts with the correct relative paths they must be executed from the ~/polhe3/production/hallc_online directory. They can be executed as 'root -b -q "analysis/elastic/el_asym_shms.cxx(11091)"'.
    • Output: The analysis plots can be found in ~/polhe3/production/hallc_replay/MON_OUTPUT/ANALYSIS_HISTOS/run#_analyzed/.
  • Elastic Asymmetry Script:
    • Location: /home/cdaq/polhe3/production/hallc_online/analysis/elastic/asym_elastic.py
    • Description: This script will calculate the elastic asymmetry for runs listed in /home/cdaq/polhe3/production/hallc_online/database/(S)HMS_run_table.csv. The momentum and angle arguments will display only runs with settings near those given. The elastic counts used for this calculation come from calling the el_asym_(s)hms.cxx scripts mentioned above. Acceptance cuts for elastic electrons can be changed in those scripts.
    • Output: The plot of the asymmetries is written to ~/polhe3/production/hallc_online/progress/asym_elastic_(s)hms.png.
      • Usage: 'python3 asym_elastic.py SHMS -p 2.17 -a 8.5 -t 0.1 --estimate=0.026 --factor-qe=0.80 --date-start=2020-01-12'
Positional Arguments:
arm Spectrometer Arm (HMS or SHMS)
Optional Arguments:
-h, --help show this help message and exit
-p MOMENTUM momentum setting (GeV)
-a ANGLE angle setting (degree)
-t TOLERANCE absolute tolerance for settings
--date-start START start date
--date-end END end date
--counts-min CMIN minimum event counts for valid elastic runs
--estimate ESTIMATE estimated asymmetry for this setting (without any corrections)
--factor-qe F_QE dilution factor from Quasi-Elastic asymmetries
--run-script run count script if output does not exist
  • Paddle Optimization Script:
    • Location: /home/cdaq/polhe3/production/hallc_online/analysis/hodo/draw_paddles_(s)hms.cxx
    • Usage: From /home/cdaq/polhe3/production/hallc_online/ execute 'root -b -q "analysis/hodo/draw_paddles_shms.cxx(9788)"'.
    • Output: The analysis plots can be found in ~/polhe3/production/hallc_online/paddle_plots/.

A1n 1-pass running notes

POST-run Calibrations/Measurements

NOTE: We'll need RadCon support and probably Survey and Alignment for several parts.

Misc Items

  • Hall is expected to go to Restricted Access around 11am--noon on Monday.
    • EPR photodiode replacement tentatively planned for Monday after lunch.
  • NOTE
    SHMS will be 'stuck' at last angle we reach during He3-Elastic program on Monday (> 8.5 degrees)
    • Will need Techs and Steve/Amy to rotate SHMS back to 8.5 degrees on Monday when RadCon Survey is completed and we are in Restricted
  • Establish a plan to keep the HB magnet on without people in the CH (Brad)
    • Got the go-ahead from Thia.
    • Discussed leaving Hall in Laser Controlled Access overnight (confirm this; maybe not worthwhile idea.. we do want to have others be able to enter the Hall in an emergency)
    • I've put an alarm on a number of critical PVs and system will send Text message alerts if they go out of bounds (trip, etc)
  • Target calibrations will likely run through to Monday 28th
    • Target group is meeting at 1:30pm on Monday to discuss plans (Bluejeans)
    • Talk to RadCon and S&A about schedule
  • Task requiring Tech Support
    • Crane work for the Compass optics table installation/removal
  • Unser Wire calibration(?) (Mack?)


Survey Work

  • Survey the Collimator positions before removal
  • Compass measurement initial setup of the table/compass and to survey the points on the screen afterwards Media:Compass_Mapping_Procedure.pdf
    • Pre-job walkthrough with Chris Gould on Thursday 24th; Survey work could happen on Monday 28th (Will have to confirm and update based on target calibration progress)

Remaining Target calibrations

Done:  See calibration spreadsheet below
 A. Monday-Tuesday
 - starting at 7 am on Monday when data taking ends:
 Tasks 1-5 will be Arun (7-8 am)/Murchhana (after 8 am)/Mingyu for PNMR)
 1) Move SHMS to 18 degrees, HB to 5.6 GeV/c. Keep target at longitudinal
    polarization 0 degree, oven at 208 degrees. Set the correction coils for
    this configuration. Move to pick-up position.
 2) After conditions stabilize, do NMR.
 3) Do PNMR, set frequency properly. DO PNMR/NMR calibration.
 4) Set convection heater to 9 V. Wait for condition stable, do PNMR/NMR.
 5) wait for ~3 hours, (before target enclosure opening for EPR
    photo-diode replacement), do PNMR/NMR. Change convection back to 7V.
 Task 6 (Junhao/JP)
 6) Replace EPR photo diode
      i) turn off lasers, Lockout laser fibers
      ii) with RADCON support, open target enclosure,
      iii) replace EPR photo-diode
      iv)) close target enclosure
      v) unlock laser fibers and turn on longitudinal lasers
 Task 7 Brad/Amy, Hall C techs.
 7) Move spectrometer to 8.5 degrees. Set HB to 2.1 GeV/c
 Task 8-19, Murchhana/Mingyu/Junhao
 8) EPR calibration for 0 degree
 9) Temperature test
 10) AFP loss measurement
 11) PNMR loss measurement
 12) Rotate to 180 degree
 13) Spin up at 180 degree
 13) EPR test
 14) Temperature test
 15) AFP loss measurement
 16) PNMR loss measurement
 17) Rotate to 270 degree, set QWPs
 18) Rotate SHMS to 11 degree, set HB to 7.5 GeV/c, set correction coils
 19) Spin up at 270 degree
 Wednesday: (Junhao/Mingyu/all)
 1) EPR at 270
 2) Temperature test at 270
 3) AFP loss at 270
 4) PNMR loss test
 5) Rotate to 90 degree
 6) Temperature test at 90
 7) AFP loss at 90
 8) PNMR study?
 9) Cross calibration between 90 and 180, ending at 180
 11) Convection speed study at 7V
 12) Rotate to 270 and Spin up at 270
 Thursday: (Junhao/Mingyu/all)
 1) Rotate SHMS to 18 degree, HB to 5.6 GeV/c
 2) AFP loss at 90
 3) EPR at 90
 4) ladder position test NMR/PNMR
 5) PNMR study?
 6) Study field effect on NMR
 7) Cross calibration between 90 and 0, ending at 0
 8) Convection speed study at 9 V
 9) Rotate to 90 and Spin up at 90
 10) Hot-spin down overnight
 Friday-Sunday: (all)
 1) Study possible reasons for polarization difference in different
 directions, different conditions?
 2) Study possible reasons for not reaching optimal max?
 3) Cold spin-down

Compass Field Measurements

Done: https://logbooks.jlab.org/entry/3853995
  • Compass measurements (should do as much as possible by Thurs Sep 24 so Murchhana can support)
    • Basic procedure: Media:Compass_Mapping_Procedure.pdf
    • Murchhana will update/create a procedure that minimizes number of Survey and Alignment setups (multiple screens, etc?)
    • Make a list of needed compass measurements (Murchhana)

Regular Field Measurements

  • Preparations:
    • Locate field probes, support jigs, etc.